OpenStudy (kylexchaos):

Medal and Fan! Don't want Answers! Just help! Cami goes scuba diving. She descends 18 feet in 3 minutes. Which expressions represent the average rate at which Cami’s position changes in feet per minute? Select all that are correct. A. −(18/3) B. −18/3<<< C.−18/−3 D. 6 <<<<< E.−6

OpenStudy (anonymous):

Not really sure about the answers, since they are formatted poorly, but "rate of position change" is basically a fancy way of saying "speed". Do you know the formula for speed?

OpenStudy (kylexchaos):

Yes I know the formula for speed

OpenStudy (kylexchaos):

Also, I updated the answers because I typed them wrong.

OpenStudy (anonymous):

So you need to calculate the speed here. You know that she is descending, so her speed must be negative (because she goes "down").

OpenStudy (mrnood):

well - your choices cannot be correct since one is positive , and ht eother is negative.

OpenStudy (kylexchaos):

Oh, that's right...

OpenStudy (kylexchaos):

So I'd say C and E, I'm not too sure about A although because the 3 is a positive.

OpenStudy (mrnood):

but on the assumption that 'down' is negative (this is only a convention - you COULD choose 'Up' is positive) then choose all the answers that are negative and equivalent to -18ft/3mins

OpenStudy (kylexchaos):

How do I get the equivalence to -18ft/3mins

OpenStudy (mrnood):

c is -18/-3 is this positive or negative? (and your last question? it's simple division)

OpenStudy (kylexchaos):

C is a positive

OpenStudy (kylexchaos):

So -18 divided by 3 = -6

OpenStudy (mrnood):

yes but 3 of the answers are equivalent to that

OpenStudy (kylexchaos):

So then my answers I have are C and E, correct?

OpenStudy (mrnood):

you just told me that C is positive

OpenStudy (kylexchaos):

Oh wait I meant B.

OpenStudy (kylexchaos):

B and E, I got my answer letters mixed up.

OpenStudy (kylexchaos):

B and E seem to have the most sense

OpenStudy (mrnood):

3 of the answers are equivalent

OpenStudy (kylexchaos):

3? Hm..

OpenStudy (kylexchaos):

A

OpenStudy (kylexchaos):

I believe I did my math correct

OpenStudy (mrnood):

-(18/3) = -1 * (18/3) = -1(6) = -6 you are correct

OpenStudy (kylexchaos):

Thank you!!!